OpenStudy (anonymous):

If the volume of a sphere is 36 pi cubic units, what is the radius

OpenStudy (anonymous):

\[V=4/3.\pi.r^3\]

OpenStudy (anonymous):

volume of sphere = 4/3 pie r^3 r^3 = Volume (sphere)*3/4 * pie

OpenStudy (anonymous):

just subs the values

OpenStudy (anonymous):

\[36=4/3*\pi*r^3\] \[r^3=36*3/4\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome....))